OpenStudy (thedj4jc):

Given the following system of equations, identify the type of system. x + y = 6 y = 3 - x consistent inconsistent equivalent When it was using graphs, I was able to pick up on it quickly, but this makes no sense to me.

OpenStudy (cggurumanjunath):

inconsistent

OpenStudy (thedj4jc):

Ok, could you explain how these type of problems work?

OpenStudy (texaschic101):

y = 3 - x .....y = -x + 3....slope is -1 and y intercept is 3 x + y = 6....y = -x + 6...slope is -1 and y intercept is 6 same slope, different y intercepts...parallel line...no solution...inconsistent (this means there is no solution) ====================== equivalent....means they have the same answer x + y = 9 x - y = 1 in this, x = 5 and y = 4 2x + 2y = 18 2x - 2y = 2 in this, x = 5 and y = 4 same solutions...they are equivalent ======================= consistent ...a system of equations with at least one solution y = x + 10 y = 2x on this one, x = 10 and y = 20...one solution or x + y = 2 2x + 2y = 4 this one has infinite solutions...no matter what number you put in for x or y, they will always equal

OpenStudy (thedj4jc):

Thank you!!! This makes much more sense!!